OpenStudy (joshmorris):

Which of the following expressions are equivalent to –3.5(2 – 3n) – 2.5n? A: –7 – 8n B: –7 + 8n C: –7 – 13n D: –7 – n(10.5 – 2.5) E: –7 + n(10.5 – 2.5)

Hayhayz (hayhayz):

use distributive law a(b+c) = ab+ac and then add any like terms in expression

OpenStudy (joshmorris):

Is it A

OpenStudy (joshmorris):

I mean B

OpenStudy (will.h):

–3.5(2 – 3n) – 2.5n Simplify using distrubtive property Distruibute -3.5 on what's inside the parentheses -3.5 * 2 = -7 -3.5 * -3n = 10.5n So now we have -7 + 10.5n - 2.5n combine alike terms -7 + 8n

OpenStudy (joshmorris):

Thanks dude